Skip to end of metadata
Go to start of metadata

went to RSW 03 to replace TRH's that were reading bad  at 40 m (missing) and 10 m 

  • found that ports were not working (2 and 6) so moved to 8 and 10 (2 -8, 6 -10) called Gary to change the configs 
  • turns out that both TRH's were pulling to much current and blowing the fuses - will check out the sensors tomorrow to determine the problem.
  • visitied RSW 05 for a problematic Metek sonic - also suspect a bad DSM because a replacement sonic and cable showed the same problem. will investigate the Metek parameters tomorrow as concerned with Viper DSM.
  • Added TRH's at TNW03 at 10m (port 3) and 2 meters (port 1)
  • noticed looking across the ridge that RNE 04 tower is now up. Will attempt to instrument that tower as time allows. Priority is to make sure existing network is operational.
  • searched for other towers being erected, but found no signs of activity at TrNW04 or TrSE 15 and 16. Not getting any updates from Inegi about tower progress, or them installing batteries at 2 sites still waiting for power (TrNW01,  or Va03)
  • TrSe08 now has power according to the tower update spread sheet but is not operational - will investigate tomorrow? 
  • assisted water vapor dial with manpower unloading equipment.
  • No labels